OpenStudy (anonymous):

In the literal equation 7z - 3x = 18 Z - 5x + 1 How do I know which one to flip?

OpenStudy (anonymous):

is the second eqn correct?mean where is the equality sign?

OpenStudy (anonymous):

What i mean by flip is in the equation x - y = 4 4y - x = 14 I flip the x - y = 4 to x = 4 + y

OpenStudy (anonymous):

Sorry equation is supposed to be 7z - 3x = 18 Z - 5x = 1

OpenStudy (anonymous):

look z has got coefficient 1, so its easy to handle with that

OpenStudy (amistre64):

you flip which ever one you want; make a choice and stick with it

OpenStudy (anonymous):

Because my Algebra instructor flipped the first one

OpenStudy (anonymous):

So it is a matter of choice, it doesn't matter which one you flip? You will get the same answer out of both?

OpenStudy (amistre64):

it doesnt matter which variable you start with, or which equation .... just do what you think is easiest for you to solve with ...

OpenStudy (anonymous):

that will lead to fractional value, a bot hard..nothing wrong with it anyways

OpenStudy (anonymous):

bit*

OpenStudy (anonymous):

She came up with 7z = 18 + 3x z + 18 + 3x / 7 / = fraction

OpenStudy (amistre64):

and that is completely usable ..... but does it have to be done that way? no

OpenStudy (anonymous):

and then it was 2(18 + 3x / 7) - 5x = 1

OpenStudy (amistre64):

since z=z=z=z.... it doesnt matter how you 'solve' for z; its gonna retain the same value thruout the process

OpenStudy (anonymous):

To avoid fractions (which I hate) I would flip the 2z - 5x = 1

OpenStudy (amistre64):

an orange ball is still an orange ball regardless of whether its in the bathtub or the dogpen...

OpenStudy (amistre64):

z = z ; so go with whats easiest for you to play with lol

OpenStudy (anonymous):

Are you saying either way I will get an equation?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

* Fraction

OpenStudy (amistre64):

you might end up using fractions; dunno ..... i havent solved it ;)

OpenStudy (anonymous):

I am going to try ad solve it! (The other way)

OpenStudy (anonymous):

*and

OpenStudy (amistre64):

fractions can be cleared tho; just multiply it all by the bottom number

OpenStudy (anonymous):

Now that is something I actually know. lol

OpenStudy (anonymous):

Thanks for the help once again!

OpenStudy (amistre64):

yw ;)

OpenStudy (anonymous):

I have 7z - 3x = 18 2z = 1 + 5x And now I dk what to do